# Enhancing LLM Reasoning with Iterative DPO: A Comprehensive Empirical Investigation

### v2

@@ Abstract
Recent advancements in post-training methodologies for large language models (LLMs) have highlighted reinforcement learning (RL) as a critical component for enhancing reasoning. However, the substantial computational costs associated with RL-based approaches have led to growing interest in alternative paradigms, such as Direct Preference Optimization (DPO). In this study, we investigate the effectiveness of DPO in facilitating self-improvement for LLMs through iterative preference-based learning. We demonstrate that a single round of DPO with coarse filtering significantly enhances mathematical reasoning performance, particularly for strong base model. Furthermore, we design an iterative enhancement framework for both the generator and the reward model (RM), enabling their mutual improvement through online interaction across multiple rounds of DPO. Finally, with simple verifiable rewards, our model DPO-VP achieves RL-level performance with significantly lower computational overhead. These findings highlight DPO as a scalable and cost-effective alternative to RL, offering a practical solution for enhancing LLM reasoning in resource-constrained situations.
@@ 1 Introduction
The field of large language models (LLMs) is evolving at an unprecedented pace (Hurst et al., 2024; Xi et al., 2025). With the release of o1 (Jaech et al., 2024) by OpenAI in September 2024, the research focus shifted from pre-training to post-training. This transition was further accelerated in January 2025, when DeepSeek-R1 (Guo et al., 2025) brought post-training methodologies to the forefront of LLM research. Then, several studies have explored reinforcement learning (RL) (Gandhi et al., 2025; Zeng et al., 2025; Cheng et al., 2025) at scale and supervised fine-tuning (SFT) (Muennighoff et al., 2025; Ye et al., 2025) with high-quality long-chain data, unlocking emergent capabilities such as self-checking, reflection, and verification in strong base models (e.g., Qwen2.5 (Yang et al., 2024a)). These advancements highlight the potential of post-training in enhancing LLM reasoning (Kumar et al., 2025).
Generally, searching self-generated data for training within post-training is referred to self-improvement (Yang et al., 2024b). At its core, this approach exploits search and learning to leverage the power of general methods as computational resources increase — resonating with the principle articulated by Sutton (2019) in The Bitter Lesson. Naturally, search and expansion can be achieved through Monte Carlo Tree Search (MCTS) and RL. MCTS-based test-time scaling and reward model (RM) training were once considered key contributors behind o1 (Zhang et al., 2024a; Cui et al., 2025). But the high inference cost and training complexity gradually pushed MCTS out of the research spotlight. In contrast, RL with verifiable rewards has been empirically shown to be more efficient and promising (Guo et al., 2025; Xie et al., 2025). However, current RL training still requires substantial computational resources for reproduction. For instance, SimpleRL (Zeng et al., 2025) trains for 1.5 days on 4×8 H100 GPUs, while PURE (Cheng et al., 2025) also demands 8 A100 GPUs. Achieving comparable or even superior self-improvement performance with lower computational resources and reduced training time is an ideal goal pursued in this field.
@@ 2.2 Training Reward Model with Online Generated Data
Furthermore, we can train the reward model using data generated by the online generator. Taking PRM as an example, we leverage external annotators (e.g., DeepSeek-V3 (Liu et al., 2024)) to label the responses sampled from the current policy $\pi_{t}$. This process produces a new round of PRM training data, which is then used to fine-tune the previous PRM $R_{t-1}$, yielding an updated PRM $R_{t}$. The training loss for updating the PRM is defined as follows:
@@ 3.1 Setup
First, we need to ensure that the base model can align its output with a specified step-by-step format. To achieve this, we select problems from the training sets of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) and reconstruct the reasoning paths in a structured step-by-step format using DeepSeek-V3 (Liu et al., 2024). The specific prompt can be found in Figure 5 in the Appendix A. This process facilitates the creation of a well-structured SFT dataset. Subsequently, we conduct two epochs of SFT on the base model.
@@ 3.5.1 Setup and Main Results
Finally, we conduct multi-round DPO on the advanced Qwen2.5-Math-7B, aiming to benchmark against RL-based approaches under the same base model and training data. Inspired by DeepSeek-R1 (Zeng et al., 2025), we observe that the verifiable reward (VR) is a 0/1 discrete variable, indicating whether the output meets predefined correctness criteria. This aligns with the original DPO framework, where correct samples are classified as positive and incorrect ones as negative, forming verifiable pairs (VP) for training. By applying VP-based filtering and gradual temperature scaling (inspired by Xi et al. (2024)), we conduct 6 epochs of DPO. The final model, Qwen2.5-7B-DPO-VP, is compared against several state-of-the-art RL-based methods:
@@
    Qwen2.5-7B-Simple-RL-Zero (Zeng et al., 2025): A model trained by RL with verifiable rewards without any extra SFT data.
@@
To align with existing RL methods, following Zeng et al. (2025) and Cheng et al. (2025), we select 8K Level 3-5 questions from the MATH as self-improvement data, and additionally evaluate on Olympiad-level benchmarks, including AMC23, AIME24, and OlympiadBench He et al. (2024). Furthermore, we conduct additional comparisons on two out-of-domain datasets to assess generalization performance. In addition, we compare with Qwen2.5-7B-DPO-R1-Zero (Zhang et al., 2025a) that applies iterative DPO on 200K questions (more than 8K questions) from the NuminaMath dataset (Li et al., 2024a). Moreover, we employ annealed sampling to enhance response diversity. Specifically, the sampling temperature $t$ is set to 0.7 for the first three epochs, increased to 1.0 during epochs 4–5, and further raised to 1.2 in the final epoch. This annealing schedule is motivated by the observation that model performance nearly converges after three epochs, and increasing the temperature in later stages allows for fine-tuning on more diverse samples, leading to additional gains.
@@
Qwen2.5-7B-Simple-RL-Zero 77.2 33.5 37.6 62.5 33.3 48.8
@@
Self-improve Data 66K $\sim$3.647M 150K 8K 8K 8K
GPUs & Training Times - 80 H100 for few days (at most) 8 A100 for few days 32 H100 for 1.5 days 8 A100 for 1 day 4 A800 for $<$1 day or 1 A800 for 3 days
@@ 3.5.2 More Analysis
First, we analyze token length variations throughout training. In Figure 4, we illustrate the accuracy progression of the iterative process across different datasets in relation to the average token length. The results indicate a steady improvement in accuracy, while the token length during inference did not exhibit an initial increase followed by a decline. Instead, it remains relatively stable within a consistent range. When comparing different models, we observe that their output lengths are generally similar. Detailed results are provided in Table 6 in Appendix E.
@@
We do not observe a significant ”Aha Moment,” which aligns with the findings of PURE (Cheng et al., 2025). Even when using CoT prompts, Qwen2.5-Math-7B frequently relies on Python-based solutions, despite not actually executing the code. Additionally, we identify a recurring self-evaluation pattern, such as ”Let’s re-evaluate …”, which is already present in the base model. This suggests that DPO does not inherently introduce self-reflection capabilities but instead strengthens existing tendencies through reward optimization, as shown in Figure 4. When comparing several different models, Table 7 in Appendix E shows that Qwen2.5-Math-7B-Instruct exhibits significantly fewer occurrences of ”Let’s”, indicating a shift in reasoning paradigms.
Evolution of Reasoning in DPO-Optimized Models
Why Iterative DPO Leads to Improvement
Based on the rigorous theoretical analysis in Xiong et al. (2024), intuitively, iterative DPO leads to improvement because it allows the model to continually refine its policy through updated feedback, thereby reducing reward misalignment over time. Additionally, the batch exploration strategy promotes diverse sampling, which enhances generalization and mitigates overfitting to early-stage preferences. The performance gains of iterative DPO on math tasks stem from the inherent self-reflection ability in the base model and the iterative nature of the near-online RL setting, as further discussed in Appendix F.
@@ 4 Related Works
Self-improvement primarily occurs during the training phase, where the model enhances its own performance by evaluating and refining its outputs without relying on external human supervision (Xi et al., 2024). It is generally believed that LLMs are incapable of achieving self-improvement without any external information (Huang et al., 2024; Tyen et al., 2024). Consequently, an increasing number of methods leverage external supervision signals help self-improvement, for example, outcome labels for mathematical or programming problems (McAleese et al., 2024; Gao et al., 2024), process reward models (Zhang et al., 2024b), value function (Wang et al., 2024a) or LLM-as-a-judge (Li et al., 2024b; Zhang et al., 2025b) frameworks. The most closely related contemporary works to ours are (Xiong et al., 2024), which proposed the iterative DPO with a rigorous theoretical analysis to demonstrate the advantage of online iterative training in the context of chat task. They further generalize the iterative DPO to the multi-turn scenario and take the potential external messages into consideration to tool-integrated reasoning tasks (Xiong et al., 2025a). In contrast, we conduct a preliminary study on data selection strategies in DPO and propose an iterative training framework for generator-verifier models, enabling mutual enhancement between the generator and the RM.
@@
In early 2025, with the release of DeepSeek-R1 (Guo et al., 2025), efficient post-training paradigms gained significant attention (Trung et al., 2024). SimpleRL-Zero (Zeng et al., 2025) achieved remarkable performance through RL, training on just 8K challenging mathematical problems with rule-based rewards. Xiong et al. (2025b) integrated reflective vocabulary into generative RM to enable selective RL-based correction. PURE (Cheng et al., 2025) revisited stepwise reward credit assignment in RL. Logic-RL (Xie et al., 2025) refined the penalty function for formatting errors, leading to improved logical reasoning capabilities. From the data utilization perspective, LIMO (Ye et al., 2025) and S1 (Muennighoff et al., 2025) demonstrated that fine-tuning with carefully curated long-chain reflective data can effectively amplify the latent reasoning capabilities. From the reasoning paradigm perspective, CFT (Wang et al., 2025) encouraged models to develop critical thinking during SFT. Our work focuses on achieving RL-based reasoning with significantly fewer computational resources. We demonstrate that, with a strong base model, coarse-grained selection is enough for performance enhancement, achieving competitive results with reduced computational overhead. Our overall training framework is more resource-efficient than the widely adopted SFT+RL pipeline, while outperforming concurrent iterative DPO methods.
@@ Appendix B Details of Processbench Evaluation
We evaluate the performance of the trained PRM on ProcessBench. Notably, in ProcessBench, the number of incorrect/correct samples in the four datasets are as follows: GSM8K: 207/193; MATH: 594/406; OlympiadBench: 661/339; OmniMath: 759/241. At this point, the traditional F1 score is not very effective for handling imbalanced datasets.
@@ Appendix C Details of Test-time Scaling
Appendix D Details of DPO and RM Training
In Section 3.2-3.4, data sampling, and evaluation of the generator are based on the OpenRLHF (Hu et al., 2024) framework. When filtering DPO data, the sample temperature is set to $0.7$. All training is performed using full parameter fine-tuning, with an SFT learning rate of $65e-$ 6, a DPO learning rate of $75e-$ 7, a maximum sequence length of $2048$, the coefficient $\beta$ of $0.1$ and a training batch size of $256$. For ORM and PRM training, we use the TRL (von Werra et al., 2020) framework with a learning rate of $61e-$ 6 and a batch size of $256$.
In Section 3.5, the DPO training follows the same configuration as above but employs annealed sampling to enhance response diversity. Specifically, for the first 3 epochs, the temperature $t$ is set to $0.7$; for epochs 4-5, $t=1.0$; and in the final epoch, $t=1.2$. More details and code can be found at https://github.com/TU2021/DPO-VP.

### v3

@@ Abstract
Recent advancements in post-training methodologies for large language models (LLMs) have highlighted reinforcement learning (RL) as a critical component for enhancing reasoning. However, the substantial computational costs associated with RL-based approaches have led to growing interest in alternative paradigms, such as Direct Preference Optimization (DPO). In this study, we investigate the effectiveness of DPO in facilitating self-improvement for LLMs through iterative preference-based learning. We demonstrate that a single round of DPO with coarse filtering significantly enhances mathematical reasoning performance, particularly for strong base model. Furthermore, we design an iterative enhancement framework for both the generator and the reward model (RM), enabling their mutual improvement through online interaction across multiple rounds of DPO. Finally, with simple verifiable rewards, our model DPO-VP achieves RL-level performance with significantly lower computational overhead. These findings highlight DPO as a scalable and cost-effective alternative to RL, offering a practical solution for enhancing LLM reasoning in resource-constrained situations. Code available at: https://github.com/TU2021/DPO-VP.
footnotetext: * Corresponding authors: Qichao Zhang and Xiangyuan Lan.
@@ 1 Introduction
The field of large language models (LLMs) is evolving at an unprecedented pace (Hurst et al., 2024; Xi et al., 2025). With the release of O1 (Jaech et al., 2024) by OpenAI in September 2024, the research focus shifted from pre-training to post-training. This transition was further accelerated in January 2025, when DeepSeek-R1 (Guo et al., 2025) brought post-training methodologies to the forefront of LLM research. Then, several studies have explored reinforcement learning (RL) (Gandhi et al., 2025; Zeng et al., 2025b; Tu et al., 2025b) at scale and supervised fine-tuning (SFT) (Muennighoff et al., 2025; Ye et al., 2025; Fu et al., 2025a) with high-quality long-chain data, unlocking emergent capabilities such as self-checking, reflection, and verification in strong base models (e.g., Qwen2.5 (Yang et al., 2024a)). These advancements highlight the potential of post-training in enhancing LLM reasoning.
Generally, searching self-generated data for training within post-training is referred to self-improvement (Yang et al., 2024b). At its core, this approach exploits search and learning to leverage the power of general methods as computational resources increase — resonating with the principle articulated by Sutton (2019) in The Bitter Lesson. Naturally, search and expansion can be achieved through Monte Carlo Tree Search (MCTS) and RL. MCTS-based test-time scaling and reward model (RM) training were once considered key contributors behind O1 (Zhang et al., 2024a; Cui et al., 2025). But the high inference cost and training complexity gradually pushed MCTS out of the research spotlight. In contrast, RL with verifiable rewards has been empirically shown to be more efficient and promising (Guo et al., 2025; Xie et al., 2025). However, current RL training still requires substantial computational resources for reproduction. For instance, SimpleRL (Zeng et al., 2025b) trains for 1.5 days on 4×8 H100 GPUs, while PURE (Cheng et al., 2025) also demands 8 A100 GPUs. Achieving comparable or even superior self-improvement performance with lower computational resources and reduced training time is an ideal goal pursued in this field.
@@ 2.2 Training Reward Model with Online Generated Data
Furthermore, we can train the reward model using data generated by the online generator. Taking PRM as an example, we leverage external annotators (e.g., DeepSeek-V3 (Liu et al., 2024a)) to label the responses sampled from the current policy $\pi_{t}$. This process produces a new round of PRM training data, which is then used to fine-tune the previous PRM $R_{t-1}$, yielding an updated PRM $R_{t}$. The training loss for updating the PRM is defined as follows:
@@ 3.1 Setup
First, we need to ensure that the base model can align its output with a specified step-by-step format. To achieve this, we select problems from the training sets of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) and reconstruct the reasoning paths in a structured step-by-step format using DeepSeek-V3 (Liu et al., 2024a). The specific prompt can be found in Figure 6 in the Appendix A. This process facilitates the creation of a well-structured SFT dataset. Subsequently, we conduct two epochs of SFT on the base model.
@@ 3.5.1 Setup and Main Results
Finally, we conduct multi-round DPO on the advanced Qwen2.5-Math-7B, aiming to benchmark against RL-based approaches under the same base model and training data. Inspired by DeepSeek-R1 (Zeng et al., 2025b), we observe that the verifiable reward (VR) is a 0/1 discrete variable, indicating whether the output meets predefined correctness criteria. This aligns with the original DPO framework, where correct samples are classified as positive and incorrect ones as negative, forming verifiable pairs (VP) for training. By applying VP-based filtering and gradual temperature scaling (inspired by Xi et al. (2024)), we conduct 6 epochs of DPO. The final model, Qwen2.5-7B-DPO-VP, is compared against several state-of-the-art RL-based methods:
@@
    Qwen2.5-7B-GRPO-Zero (Shao et al., 2024): A model trained by GRPO with verifiable rewards without any extra SFT data.
@@
    Qwen2.5-7B-Simple-RL-Zero/Zoo (Zeng et al., 2025a; b): A model trained by PPO with verifiable rewards without any extra SFT data. Zero and Zoo refer to two versions released at different time points.
  6. (f)
@@
To align with existing RL methods, following Zeng et al. (2025b) and Cheng et al. (2025), we select 8K Level 3-5 questions from the MATH as self-improvement data, and additionally evaluate on Olympiad-level benchmarks, including AMC23, AIME24, and OlympiadBench He et al. (2024). Furthermore, we conduct additional comparisons on 5 out-of-domain datasets to assess generalization performance. In addition, we compare with Qwen2.5-7B-DPO-R1-Zero (Zhang et al., 2025a) that applies iterative DPO on 200K questions (more than 8K questions) from the NuminaMath dataset (Li et al., 2024a).
Moreover, we employ annealed sampling to enhance response diversity. Specifically, the sampling temperature $t$ is set to 0.7 for the first three epochs, increased to 1.0 during epochs 4–5, and further raised to 1.2 in the final epoch. This annealing schedule is motivated by the observation that model performance nearly converges after three epochs, and increasing the temperature in later stages allows for fine-tuning on more diverse samples, leading to additional gains.
@@
Qwen2.5-7B-GRPO-Zero 76.2 32.7 38.1 55.0 16.7 43.7
Qwen2.5-7B-Simple-RL-Zero 78.0 33.1 36.6 60.0 26.7 46.9
Qwen2.5-7B-Simple-RL-Zoo 80.4 39.7 38.8 57.5 26.7 48.7
@@
Self-improve Data 66K $\sim$3.647M 150K 8K MATH 8K MATH 8K MATH
GPUs & Training Times - 80 H100 for few days (at most) 8 A100 for few days 32 H100 for 1.5 days 8 A100 for 1 day 4 A800 for $<$1 day or 1 A800 for 2 days
@@ 3.5.2 More Analysis
First, we analyze token length variations throughout training. In Figure 4, we illustrate the accuracy progression of the iterative process across different datasets in relation to the average token length. The results indicate a steady improvement in accuracy, while the token length during inference did not exhibit an initial increase followed by a decline. Instead, it remains relatively stable within a consistent range. When comparing different models, we observe that their output lengths are generally similar. Detailed results are provided in Table 7 in Appendix E.1.
@@
We do not observe a significant ”Aha Moment,” which aligns with the findings of PURE (Cheng et al., 2025). Even when using CoT prompts, Qwen2.5-Math-7B frequently relies on Python-based solutions, despite not actually executing the code. Additionally, we identify a recurring self-evaluation pattern, such as ”Let’s re-evaluate …”, which is already present in the base model. This suggests that DPO does not inherently introduce self-reflection capabilities but instead strengthens existing tendencies through reward optimization, as shown in Figure 4. When comparing several different models, Table 8 in Appendix E.1 shows that Qwen2.5-Math-7B-Instruct exhibits significantly fewer occurrences of ”Let’s”, indicating a shift in reasoning paradigms.
3.6 Ablation Study on the DPO Sampling Noisy
Figure 5: Label Noise Ablation
To evaluate the robustness of our DPO framework under noisy supervision, we conduct a controlled label noise ablation. Specifically, we simulate noise in the preference data used to train the Preference Reward Model (PRM) under the PRM with offset setting. Following Tu et al. (2025c), for each positive–negative response pair, we apply a random flipping probability $p\in[0,1]$ that reverses the preference direction with probability $p$, simulating increasing levels of annotation noise. For example, $p=0.5$ corresponds to fully random pairwise supervision, while $p=1.0$ denotes complete inversion of the labels.
We test the impact of this noise injection on the final model performance, measured by pass@1 accuracy after one round of DPO on GSM8K and MATH500 using Qwen2.5-7B. The results are shown in Figure 5. These results clearly demonstrate that model performance degrades as the preference supervision becomes less accurate. While DPO appears tolerant to small levels of label noise, performance drops significantly under higher noise levels. This provides empirical evidence that maintaining reasonably accurate verifier signals is critical for effective preference optimization.
Generalization
We conduct evaluations on 5 additional Out-of-Domain datasets:
  1. (a)
    CMATH: A Chinese mathematics dataset.
  2. (b)
    MMLU_STEM: A subset of MMLU focusing on science subjects like physics, biology.
  3. (c)
    HumanEval: A benchmark evaluating code generation through Python tasks.
  4. (d)
    LiveCodeBench: A coding benchmark for real-time code editing and debugging.
  5. (e)
    RACE: A reading benchmark from English exams testing reasoning.
The first two datasets are evaluated using the Qwen-Math codebase (Yang et al., 2024b), with evaluation parameters consistent with those used in Appendix D. The remaining three datasets are evaluated via EvalScopehttps://github.com/modelscope/evalscope using its default configuration settings. The results are summarized in Table 4.
pass@1 acc CMATH MMLU_STEM HumanEval Live_Code_Bnech RACE Average
Qwen2.5-Math-7B 64.0 58.3 40.9 5.1 61.6 46.0
Eurus-2-7B-PRIME 72.5 42.8 41.5 21.2 62.6 48.1
Qwen2.5-7B-Simple-RL-Zero 70.3 59.1 51.8 21.4 61.1 52.7
Qwen2.5-7B-Simple-RL-Zoo 89.0 35.9 53.1 17.9 63.9 52.0
Qwen2.5-7B-PURE-VR 69.5 58.8 43.3 6.4 61.3 47.9
Qwen2.5-7B-DPO-VP 68.5 59.3 50.6 17.3 61.2 51.4
Table 4: Performance comparison on Out-of-Domain datasets.
Nearly all methods lead to slight performance improvements on the out-of-domain (OOD) benchmarks, indicating enhanced generalization capabilities. Our proposed DPO-VP notably increases the average OOD score from 46.0 to 51.4, achieving gains comparable to those of RL-based methods. This suggests that DPO fine-tuning on mathematical tasks can also effectively enhance generalization to other domains.
@@ 4 Related Works
Self-improvement primarily occurs during the training phase, where the model enhances its own performance by evaluating and refining its outputs without relying on external human supervision (Xi et al., 2024; Kumar et al., 2025). It is generally believed that LLMs are incapable of achieving self-improvement without any external information (Huang et al., 2024; Tyen et al., 2024). Consequently, an increasing number of methods leverage external supervision signals help self-improvement, for example, outcome labels for mathematical or programming problems (McAleese et al., 2024; Gao et al., 2024), process reward models (Zhang et al., 2024b; Fu et al., 2025b), value function (Wang et al., 2024a) or LLM-as-a-judge (Li et al., 2024b; Zhang et al., 2025b) frameworks.
Improving Reasoning with DPO
Recent works have significantly extended DPO-based reasoning (Xu et al., 2024; Zhang et al., 2025a). Iterative DPO with rigorous theoretical analysis was proposed by Xiong et al. (2024), and further generalized to multi-turn and tool-integrated scenarios (Xiong et al., 2025a). Flow-DPO (Deng & Mineiro, 2024) introduced online multi-agent learning for mathematical reasoning, while iterative length-regularized DPO (Liu et al., 2024b) enables 7B models to approach GPT-4-level reasoning. Self-play and self-training pipelines have also been explored, demonstrating that DPO and preference-based methods can improve LLMs beyond human-labeled data (Chen et al., 2024b; Singh et al., 2023). Additionally, implicit reward bootstrapping with DPO has shown promise for further alignment without explicit reward models (Chen et al., 2024a). Aya Expanse (Dang et al., 2024) illustrates how such alignment advances benefit multilingual and cross-cultural reasoning. In contrast, our work focuses on data selection strategies for DPO and proposes an iterative training framework for generator-verifier models, enabling mutual enhancement between generation and reward modeling.
@@
In early 2025, with the release of DeepSeek-R1 (Guo et al., 2025), efficient post-training paradigms gained significant attention (Trung et al., 2024; Tu et al., 2025a). SimpleRL-Zero (Zeng et al., 2025b) achieved remarkable performance through RL, training on just 8K challenging mathematical problems with rule-based rewards. Xiong et al. (2025b) integrated reflective vocabulary into generative RM to enable selective RL-based correction. PURE (Cheng et al., 2025) revisited stepwise reward credit assignment in RL. Logic-RL (Xie et al., 2025) refined the penalty function for formatting errors, leading to improved logical reasoning capabilities. From the data utilization perspective, LIMO (Ye et al., 2025) and S1 (Muennighoff et al., 2025) demonstrated that fine-tuning with carefully curated long-chain reflective data can effectively amplify the latent reasoning capabilities. From the reasoning paradigm perspective, CFT (Wang et al., 2025) encouraged models to develop critical thinking during SFT. Our work focuses on achieving RL-based reasoning with significantly fewer computational resources. We demonstrate that, with a strong base model, coarse-grained selection is enough for performance enhancement, achieving competitive results with reduced computational overhead. Our overall training framework is more resource-efficient than the widely adopted SFT+RL pipeline, while outperforming concurrent iterative DPO methods.
@@ 5 Conclusion
Acknowledgements
This work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences under Grant XDA0480303, Young Scientists Fund of The State Key Laboratory of Multimodal Artificial Intelligence Systems ES2P100112, Beijing Natural Science Foundation - Xiaomi Innovation Joint Fund L253007, and National Natural Science Foundation of China 62402252.
@@ Appendix B Details of Processbench Evaluation
We evaluate the performance of the trained PRM on ProcessBenchhttps://github.com/QwenLM/ProcessBench. Notably, in ProcessBench, the number of incorrect/correct samples in the four datasets are as follows: GSM8K: 207/193; MATH: 594/406; OlympiadBench: 661/339; OmniMath: 759/241. At this point, the traditional F1 score is not very effective for handling imbalanced datasets.
@@ Appendix C Details of Test-time Scaling
Appendix D Details of DPO and RM Training and Evaluation
[…]
In Section 3.5, the DPO training follows the same configuration as above but employs annealed sampling to enhance response diversity. Specifically, for the first 3 epochs, the temperature $t$ is set to 0.70.70.7; for epochs 4-5, $t=1.0$; and in the final epoch, $t=1.2$.
All evaluations are conducted using the Qwen-Math evaluation codebasehttps://github.com/QwenLM/Qwen2.5-Math, with greedy decoding of $t=0.0$, generating one output per input, and a maximum generation length of 204820482048 tokens. In addition, we provide the URLs of all self-evaluated models in footnotes, including Eurus-2-7B-PRIMEhttps://huggingface.co/PRIME-RL/Eurus-2-7B-PRIME, Qwen2.5-7B-Simple-RL-Zerohttps://huggingface.co/hkust-nlp/Qwen-2.5-Math-7B-SimpleRL-Zero, Qwen2.5-7B-Simple-RL-Zoohttps://huggingface.co/hkust-nlp/Qwen-2.5-Math-7B-SimpleRL-Zoo, and Qwen2.5-7B-PURE-VRhttps://huggingface.co/jinachris/Qwen2.5-7B-PURE-VR.
More details and code can be found at https://github.com/TU2021/DPO-VP.